A rifle is fired straight up, and the bullet leaves the rifle with an initial velocity magnitude of 724 m/s. After 5.00 s the velocity is 675 m/s. At what rate is the bullet decelerated?
Answer #1

We have Newton’s motion formula, v = u + at

where v is final velocity,

u is initial velocity,

a is acceleration and

t is time

thus for above question, 675 = 724 + a*5

   ==> 5a = 675-724

==>5a = -49

Therefore, a = -49/5 = -9.8 m/s2[Answer]

Negative sign indicates retardation. If magnitude is asked in answer just put 9.8 m/s2 and if both magnitude and direction is asked put -9.8 m/s2.

Note: the value we got is value of acceleration due to gravity and which is true.